# The Challenges of Evaluating LLM Applications: An Analysis of Automated, Human, and LLM-Based Approaches

### v1

@@ 5. Discussion
The goal of this work is to illustrate how challenging it is to evaluate an LLM based application, especially evaluating a chatbot with current methodologies including automated, human and LLM procedures. We first demonstrate that there are advantages and disadvantages in all three of these approaches. We also note the differences of results gained from all three evaluation procedures, there is very little correlation between these results and it would be difficult to suggest one to be used. We also observed that the experts evaluation results are a bit stricter and resulted lower scores generally for many factors. The novice had looked at the chatbot in a favorable light and we notice the slightly elevated scores. Using an LLM to evaluate the chatbot responses seems to be not reliable as the LLM scores its own responses high. In our experimental case, we used the same LLM (GPT-3.5) to generate the responses and also as the evaluator LLM. This is not the ideal setting as (svikhnushina_approximating_2023) points out, in (svikhnushina_approximating_2023) authors point out if an LLM is not evaluated it must be evaluated using an already evaluated LLM or a higher order LLM. Given this situation of uncertain evaluations from any procedure, we should not distract the readers from the need for evaluating. To improve the reliability of evaluation, we suggest increasing the number of humans used in the factored human evaluation. Also enlisting a wide range of expertise would create a smoothed preview of the results; however, this would increase the expensiveness of the evaluation. As (van_der_lee_best_2019) suggests, enlisting a larger amount of intended users of a chatbot would still not be ideal as these users may also create confusion on whats correct and whats not. Allowing untrained humans to make judgments on the factors will not yield the most accurate results, similar to the case we have with LLM results in Figure 1.
@@
One deciding factor would be the repeatability and the amount of funds a person has toward evaluating a chatbot. In this regard we note while automated procedures are repeatable, low reliability of these metrics make a case against them. Human evaluation is considered the gold standard, while that can be true research indicates that the human disagreement is a greater issue; we also notice this issue indicated in Table 6. LLM evaluators are a novel adaptation of LLMs, its greatest adversary right now is not having enough research to support its reliability. We observe that in some cases LLM evaluators have similar responses to human evaluators. But this is not the case always, in most instances LLM evaluators tend to be overly confident in the response being correct. We cannot reject the promise in LLM evaluators as we can set various personalities and take various versions of its evaluation rapidly (chan_chateval_2023), but this also must be explored in terms of whether a person of such an expertise would rate the same response in a similar way. Further research needs to be conducted in understanding how LLMs can help us evaluate LLMs.

### v2

@@ 5 Discussion
The goal of this work is to illustrate how challenging it is to evaluate an LLM based application, especially evaluating a chatbot with current methodologies including automated, human and LLM procedures. We first demonstrate that there are advantages and disadvantages in all three of these approaches. We also note the differences of results gained from all three evaluation procedures, there is very little correlation between these results and it would be difficult to suggest one to be used. We also observed that the experts evaluation results are a bit stricter and resulted lower scores generally for many factors. The novice had looked at the chatbot in a favorable light and we notice the slightly elevated scores. Using an LLM to evaluate the chatbot responses seems to be not reliable as the LLM scores its own responses high. In our experimental case, we used the same LLM (GPT-3.5) to generate the responses and also as the evaluator LLM. This is not the ideal setting as [24] points out, in [24] authors point out if an LLM is not evaluated it must be evaluated using an already evaluated LLM or a higher order LLM. Given this situation of uncertain evaluations from any procedure, we should not distract the readers from the need for evaluating. To improve the reliability of evaluation, we suggest increasing the number of humans used in the factored human evaluation. Also enlisting a wide range of expertise would create a smoothed preview of the results; however, this would increase the expensiveness of the evaluation. As [13] suggests, enlisting a larger amount of intended users of a chatbot would still not be ideal as these users may also create confusion on whats correct and whats not. Allowing untrained humans to make judgments on the factors will not yield the most accurate results, similar to the case we have with LLM results in Figure 1.
@@
One deciding factor would be the repeatability and the amount of funds a person has toward evaluating a chatbot. In this regard we note while automated procedures are repeatable, low reliability of these metrics make a case against them. Human evaluation is considered the gold standard, while that can be true research indicates that the human disagreement is a greater issue; we also notice this issue indicated in Table 6. LLM evaluators are a novel adaptation of LLMs, its greatest adversary right now is not having enough research to support its reliability. We observe that in some cases LLM evaluators have similar responses to human evaluators. But this is not the case always, in most instances LLM evaluators tend to be overly confident in the response being correct. We cannot reject the promise in LLM evaluators as we can set various personalities and take various versions of its evaluation rapidly [21], but this also must be explored in terms of whether a person of such an expertise would rate the same response in a similar way. Further research needs to be conducted in understanding how LLMs can help us evaluate LLMs.
@@ Acknowledgements.
Appendix A Prompts
This section notes the prompts that have been used in this work, we first note the prompt that has been utilized in the RAG process in the chatbot for clarity and then a sample prompt that was
A.1 RAG Prompt
The user asks the question <question>. Here are some facts that could be used to support the question, <facts delimited by semicolons>.
You must first investigate if it is possible to support an answer with the available facts If you do not have facts to support an answer, step by step explaining your reasoning behind each action you must come up with a answer by processing, applying and evaluating facts as needed. Otherwise you must only respond with "I dont know" and do not output anything else.
A.2 LLM Evaluator Prompt
Here in this prompt we only add the prompt used with the “Correctness” criterion and similar prompts can be drawn for others.
You are an expert education researcher.
You are given a set of facts, a question that relates to the text of these facts and an answer for the given question.
Your task is to evaluate if the answer is a good answer to the given question based off of a criterion and also considering the facts.
Evaluation steps:
1. Read the facts: Start by carefully reading the facts provided. Understand the context, main points, and any relevant details.
2. Analyze the Question: Examine the question that relates to the facts. Ensure you have a clear understanding of what the question is asking for.
3. Review the Answer: Carefully read the answer provided and assess it based only on the following criterion:
Correctness: Does the answer provide accurate information based on the paragraph text?
4. Assign a Score: Use the 5-point scale to assign a score to the answer:
Score 1: If the answer is wrong compared with the facts for the question
Score 2: If the answer is mostly wrong compared with the facts for the question
Score 3: If the answer is partly correct given the facts for the question
Score 4: If the answer is mostly correct given the facts for the question
Score 5: If the answer is correct given the facts for the question
5. Document Scores: Keep a record of the scores and feedback for reference.
This can be helpful for tracking progress and ensuring consistency in your evaluations.